A gas in a large sealed gas syringe absorbs 4Kj of heat thi causes the gas to expand reversibly and forces the piston with surface area 80cm² move 200cm upward against an external pressure of 690 mmHg calculate the change in internal energy
1
Expert's answer
2020-11-22T11:26:39-0500

p = 690 mmHg = 91992 Pa

A(surface area) = 80cm² = 80/(100*100) = 0.008 m²

l = 200cm = 2 m

W (work) = p*A*l = 91992*0.008*2 = 1472 J

Q = 4kJ = 4000 J

"\Delta"U (change in internal energy) = Q-W = 4000-1472 = 2528 J